# Bootstrap3D: Improving 3D Content Creation with Synthetic Data

### v1

@@ 3.3 Training Timestep Reschedule (TTR)
Despite retaining only relatively high-quality synthetic data with minimal motion blur from SV3D [76] through MV-LLaVA, small areas of blurring persist, stemming from both motion and out-of-distribution scenarios for SV3D and SVD [6]. These blurred data can potentially compromise the final performance of the multi-view diffusion model. To restrict the training time step for synthetic data, we proposed a simple yet effective Training Timestep Reschedule (TTR) method.
Background. Before delving into TTR, we briefly review some basic concepts needed to understand diffusion models (DDPMs) [30, 68, 63, 62, 12]. Gaussian diffusion models assume a forward noising process which gradually applies noise to real data $x_{0}$
@@
During training, $t$ is randomly sampled in $[0,N]$ ($N=1000$ in  [12, 62]) for the model to predict the added noise $\epsilon_{t}$, where $x_{0}$ denotes for the clear nature image and $x_{N}$ denotes for pure random noise. As depicted in Fig.3, when $t$ is large, the denoising process primarily focuses on determining the global low frequency($f$) content such as overall structure and shape. Conversely, when $t$ is small, the denoising process is predominantly responsible for generating high $f$ components such as texture.
Simple implementation. When adapting Stable Diffision [62] for multi-view generation, the previous approach [66] changes the default scaled linear schedule into the linear schedule to emphasize more on early denoising stage for structural variation and view consistency. Inspired by this, we propose restricting the denoising time step of synthetic data during training. As small yet observable blur still exists in synthetic data with novel view generated by SV3D [76], we limit them to training diffusion model only with large $t$. This restricts the backpropagation of these synthetic data to focus on the low-frequency component of the image like the overall structure and shape that faithfully follow text prompts and consistency between different views. Small $t$ values are only sampled on clear and physically consistent multi-view images rendered from Objaverse [20] and supplemented high-quality 2D images from SA-1B [38], help model outcome high-quality images with more photo-realistic and varied texture details.
@@ 4.1 Experiment Settings
Training data. For each set of 4-view images depicting the same object, obtained from both Objaverse [20] and generated by SV3D [76] or Zero123++[66], we use MV-LLaVA to generate long descriptive captions with predicted quality score. Detailed quality check of MV-LLaVA is supplied in Sup. C and data analysis in Sup. B. In the end, we generate 200K 4-view image-text pairs on Objaverse [20], 1000K 4-view image-text pairs from synthetic data from SV3D [76] and Zero123++[66]. We also sample 35K HQ SA [38] data with captions from ShareGPT4V [13].
@@
Training details. We test our framework directly on the text-to-multi-view diffusion model. We fine-tune PixArt-$\alpha$ [12] with backbone DiT-XL/2 [57] model on the data as mentioned earlier. Similar to Instant3D [41], we train the diffusion model directly on 4-view images naturally arranged in a 2$\times$2 grid. For 4 same view images from SA [38], we limit training time step $t\in[0,50]$. We limit synthetic multi-view images $t\in[200,1000]$. Regarding 3D object-rendered images, we do not limit $t$ but sample more frequently in the range $[50,200]$ as a complement. We set the total batch size to 1024 with the learning rate set to 8e-5 for 20K steps. Training is conducted on 32 NVIDIA A100-80G GPUs for 20 hours with Flan-T5-XXL [17] text features and VAE [37] features pre-extracted.
@@
Evaluation metrics. We primarily benchmark the quantitative results of our approach and other methods from two main dimensions: 1). Image-text alignment measured by CLIP score and CLIP-R score indicating the prompt follow ability of text-to-multi-view (T2MV) diffusion model. 2). Quality of generated images measured by FID [29]. Given the trend of decoupling multi-view image generation and sparse view reconstruction, we conduct tests separately on multi-view images by T2MV and rerendered images from generated 3D objects. To test the robustness and diversity of Bootstrap3D beyond prompts generated by GPT, we also collect real user prompts from public website, the details and test results are available in Sup. A.
@@
Evaluation details. For CLIP-R Score and CLIP Score, we test on 110 text prompts from GPTeval3D [85] using different CLIP models (i.e., CLIP-L/14 [61] and CLIP-bigG [34]) following the same setting of Instant3D [41]. Regarding the FID [29] test, as there is no golden standard for HQ 3D objects, we follow the similar evaluation idea of PlayGround2.5 [40] (PG2.5) to use powerful T2I model generated images to form ground truth (GT) distribution. We use curated prompts to guide powerful PixArt and PG2.5 to generate high-quality CAD-style images with a single object in the pure background. Rembg [25] is adopted to create white background object-centric images. We use the method proposed in GPTeval3D [85] to generate 3K prompts. For both PG-2.5 and PixArt, we generate 10 images for each prompt with different seeds, resulting in 30K images to form the GT distribution of high-quality CAD-style objects.
@@
Comparing methods. In addition to Instant3D [41] and MVDream [67] as direct text-to-multi-view (T2MV) methods, we also adopt edge-cutting single image to multi-view (I2MV) methods SV3D [76] and Zero123++[66]. For these methods, we condition the diffusion model on the single view image generated by PixArt (prompted to generate CAD-style single object-centric image). The result of the CLIP score is 3 times averaged with different seeds. For FID, we use 3 different seeds for each of the 3K prompts to generate 9K images to test the distance with GT high-quality images.
@@ 4.2 Evaluation of Multi-view Images
Table 1: Benchmark of CLIP and FID score of text-to-multi-view (T2MV) models on generated 4 view images, CLIP score tests on 110 text prompts from GPTeval3D [85] while FID [29] is measured with the distribution of 30K object-centric images generated by SOTA T2I models. For text-to-image-to-multi-view(T2I2MV), we input I2MV models with single view images generated by Pixart-$\alpha$, which superior single view image score is marked in green.
Domain Method CLIP-R Score $\uparrow$ CLIP Score $\uparrow$ FID [29] $\downarrow$
@@
T2I PixArt-$\alpha$ [12] 96.1 94.7 25.9 41.5 20.7 5.4
T2I2MV SV3D [76] 78.8 81.3 24.7 37.3 55.7 54.2
Zero123++ [66] 78.0 84.5 24.2 36.9 53.2 49.3
T2MV Instant3D [41] (unofficial) 83.6 91.1 25.6 39.2 83.2 77.9
MVDream [67] 84.8 89.3 25.5 38.4 60.2 59.2
@@
As illustrated in Tab.1, in addition to 4 view images generated by the multi-view diffusion model. Compared to other methods, the T2MV diffusion model trained by our framework yields the best results both according to image-text alignment and image quality. It is worth noting that in the T2MV task, T2I2MV suffers a huge drop due to the domain gap between T2I-generated images and training images as well as the use of Rembg [25]. Qualitative results of real user cases are available in Sup. A.
Figure 4: Bootstrap3D generates 3D objects compared to other edge-cutting methods given text prompt. More results with higher resolution are available in Sup.F.1.
@@ 4.3 Evaluation of Generated 3D Objects
Table 2: Benchmark of CLIP and FID score of generated 3D objects based on rendered 9 view images. *While MVDream is tested on 200 generated objects for FID test using SDS [67], other methods are tested on 1000 objects using GRM [89] as sparse view reconstruction model.
Method CLIP-R Score $\uparrow$ CLIP Score $\uparrow$ FID [29] $\downarrow$
CLIP-L/14 CLIP-bigG CLIP-L/14 CLIP-bigG PG2.5 [40] PixArt [12]
Instant3D [41] (unofficial) 81.7 89.4 24.8 37.1 85.4 80.3
MVDream [67]* 85.2 90.8 26.1 39.4 57.8 56.7
SV3D [76] 74.1 82.8 23.4 34.1 68.4 69.1
Zero123++ [66] 71.2 80.3 22.3 34.5 69.3 72.4
@@
Figure 5: Bootstrap3D can generate 3D objects with different style with more precise prompt control compared to other edge-cutting methods. More results with thorough render views are available in Sup.F.2.
View consistency is another crucial factor in reconstructing reasonable 3D objects. Miss alignment between different views can lead to blurred areas in reconstructed objects by large reconstruction model [32, 41, 79, 88, 84, 89]. To assess the final quality of the generated 3D object, we employ GRM [89] to reconstruct the object given sparse view images generated in Sec. 4.2. We render 9 view images evenly in orbit for each object and evaluate the image-text alignment and image quality. As reported in Tab. 2. Bootstrap3D, after conditioning GRM on 4 view images, can generate the best 3D objects both according to image-text alignment and image quality. GPT-4V based human-aligned evaluation based on GPTeval3D [85] is supplied in Sup. E.
We also present visualizations of some results in Fig.4. Compared with GRM [89], Bootstrap3D can generate objects with more diversity and prompt follow ability. Zero123++ [66] here uses the text-to-image-to-multi-view pipeline, with the first single view image generated by PixArt [12]. As shown in the first column of Fig.4, Although the first image may be well aligned with the given text prompt. The final 3D object may be compromised due to the limitations of its multi-view diffusion model as it is also fine-tuned on Objaverse [20] only. The domain gap between the rendered Objaverse image and images generated by the T2I model limits the outcome of the text-to-single image, single-view-to-3D pipeline.
@@ 4.4 Ablation Study
Table 3: Ablation study of proposed components and quantity of synthetic data. with CLIP-R Score represents image-text alignment and FID [29] represents image quality.
@@
Synthetic data and dense captioning are proposed in our work to achieve high-quality images and better image-text alignment. We ablate their effects and the importance of data quantity in Tab. 3. Direct use of synthetic data without Training Timestep Reschedule (TTR) can cause severe blurs and deformation in final outcome. With the help of TTR, the mixture of data can not only improve image-text alignment but also maintain view consistency. Replacing Cap3D [52]’s caption with MV-LLaVA’s dense descriptive caption further improves the model’s capability of following prompts faithfully. Improvement through increasing volume of data also proves the scalability of our framework.
@@ 5 Conclusion and Discussion
Limitations and future work. Despite its promise, our work still faces several unresolved challenges. Firstly, the multi-view diffusion model is only the first step of the 3D content creation pipeline. Sparse view reconstruction models also need improvement as most edge-cutting sparse view reconstruction models are also trained on Objaverse [20] only. Secondly, Although MLLMs can estimate general quality and view consistency, subtle view inconsistency is hard to detect until ambiguity leads to blurred areas in reconstructed 3D content. While the proposed Training Timestep Reschedule can mitigate this problem, it cannot solve the problem fundamentally. Using synthetic data to train sparse view reconstruction models and quality estimation directly based on the reconstructed object are thus interesting future directions for improving 3D content creation.
Appendix A Evaluation on wild prompts from real users
The results of the main part of the paper are only tested on GPT generated prompts. To test our work’s capability in wild cases, we also collect real user prompts and compare our method with Instant3D [41]. specifically, we randomly collect 100 prompts from https://www.meshy.ai/ and test the CLIP-R precision as well as GPT based evaluation (detailed in Sup. E). Results and some qualitative cases are shown in Tab. 4 and Fig. 7. We highlight that our Bootstrap3D excels Instant3D [41] when tested on real user prompts through training on synthetic data.
Table 4: Test results of in the wild cases. Bootstrap3D also excels Instant3D [41] in generating high quality images according to real user prompts.
@@ Appendix A Evaluation on wild prompts from real users
Figure 7: Real user prompt cases visualization compared to Instant3D [41]
Appendix B Data Statistics
B.1 Caption Analysis
Figure 8: Visualized analysis of dense reasoning descriptions generated by GPT4-Vision [55] of the root noun-verb pairs (occurring over 1%) of the descriptions

### v2

@@ 3.3 Training Timestep Reschedule (TTR)
Despite retaining only relatively high-quality synthetic data with minimal motion blur from SV3D (Voleti et al., 2024) through MV-LLaVA, small areas of blurring persist, stemming from both motion and out-of-distribution scenarios for SV3D and SVD (Blattmann et al., 2023). These blurred data can potentially compromise the final performance of the multi-view diffusion model. To restrict the training time step for synthetic data, we proposed a simple yet effective Training Timestep Reschedule (TTR) method.
Background. Before delving into TTR, we briefly review some basic concepts needed to understand diffusion models (DDPMs) (Ho et al., 2020; Sohl-Dickstein et al., 2015; Salimans & Ho, 2022; Rombach et al., 2022; Chen et al., 2023a). Gaussian diffusion models assume a forward noising process which gradually applies noise to real data $x_{0}$
@@
During training, $t$ is randomly sampled in $[0,N]$ ($N=1000$ in  (Chen et al., 2023a; Rombach et al., 2022)) for the model to predict the added noise $\epsilon_{t}$, where $x_{0}$ denotes for the clear nature image and $x_{N}$ denotes for pure Gaussian noise. As depicted in Fig.4, when $t$ is large, the denoising process primarily focuses on determining the global low frequency($f$) content such as overall structure and shape. Conversely, when $t$ is small, the denoising process is predominantly responsible for generating high $f$ components such as texture.
When adapting Stable Diffision (Rombach et al., 2022) for multi-view generation, the previous approach (Shi et al., 2023a) changes the default scaled linear schedule into the linear schedule to emphasize more on early denoising stage for structural variation and view consistency. Inspired by this, we propose restricting the denoising time step of synthetic data during training. As small yet observable blur still exists in synthetic data with novel view generated by SV3D (Voleti et al., 2024), we limit them to training diffusion model only with large $t$. This restricts the backpropagation of these synthetic data to focus on the low-frequency component of the image like the overall structure and shape that faithfully follow text prompts and consistency between different views. Small $t$ values are only sampled on clear and physically consistent multi-view images rendered from Objaverse (Deitke et al., 2023) and supplemented high-quality 2D images from SA-1B (Kirillov et al., 2023), help model outcome high-quality images with more photo-realistic and varied texture details.
Figure 5: Bootstrap3D generates 3D objects compared to other edge-cutting methods given text prompt. More results with higher resolution are available in Sup.A.8.1.
@@ 4.1 Experiment Settings
Training data. For each set of 4-view images obtained from both Objaverse (Deitke et al., 2023) and generated by SV3D (Voleti et al., 2024) or Zero123++(Shi et al., 2023a), we use MV-LLaVA to generate long descriptive captions with predicted quality score. Detailed quality check of MV-LLaVA is supplied in Sup. A.4 and data analysis in Sup. A.3. In the end, we generate 200K 4-view image-text pairs on Objaverse (Deitke et al., 2023), 1000K 4-view image-text pairs from synthetic data from SV3D (Voleti et al., 2024) and Zero123++(Shi et al., 2023a). We also sample 35K HQ SA (Kirillov et al., 2023) data with captions from ShareGPT4V (Chen et al., 2023b).
@@
Training details. We test our framework directly on the text-to-multi-view diffusion model. We fine-tune PixArt-$\alpha$ (Chen et al., 2023a) with backbone DiT-XL/2 (Peebles & Xie, 2023) model on the data as mentioned earlier. Similar to Instant3D (Li et al., 2023a), we train the diffusion model directly on 4-view images naturally arranged in a 2$\times$2 grid. For 4 same view images from SA (Kirillov et al., 2023), we limit training time step $t\in[0,50]$. We limit synthetic multi-view images $t\in[200,1000]$. Regarding 3D object-rendered images, we do not limit $t$ but sample more frequently in the range $[50,200]$ as a complement. We set the total batch size to 1024 with the learning rate set to 8e-5 for 20K steps. Training is conducted on 32 NVIDIA A100-80G GPUs for 20 hours with Flan-T5-XXL (Chung et al., 2024) text features and VAE (Kingma & Welling, 2013) features pre-extracted.
@@
Evaluation metrics. We primarily benchmark the quantitative results of our approach and other methods from two main dimensions: 1). Image-text alignment measured by CLIP score and CLIP-R score indicating the prompt follow ability of text-to-multi-view (T2MV) diffusion model. 2). Quality of generated images measured by FID (Heusel et al., 2017). Given the trend of decoupling multi-view image generation and sparse view reconstruction, we conduct tests separately on multi-view images by T2MV and rerendered images from generated 3D objects. To test the robustness and diversity of Bootstrap3D beyond prompts generated by GPT, we also collect real user prompts from public website, the details and test results are available in Sup. A.1.
@@
Evaluation details. For CLIP-R Score and CLIP Score, we test on 110 text prompts from GPTeval3D (Wu et al., 2024) using different CLIP models (i.e., CLIP-L/14 (Radford et al., 2021) and CLIP-bigG (Ilharco et al., 2021)) following the same setting of Instant3D (Li et al., 2023a). Regarding the FID (Heusel et al., 2017) test, as there is no golden standard for HQ 3D objects, we follow the similar evaluation idea of PlayGround2.5 (Li et al., 2024) (PG2.5) to use powerful T2I model generated images to form ground truth (GT) distribution. We use curated prompts to guide powerful PixArt and PG2.5 to generate high-quality CAD-style images with a single object in the pure background. Rembg (etc, 2020) is adopted to create white background object-centric images. We use the method proposed in GPTeval3D (Wu et al., 2024) to generate 3K prompts. For both PG-2.5 and PixArt, we generate 10 images for each prompt with different seeds, resulting in 30K images to form the GT distribution of high-quality CAD-style objects.
@@
Comparing methods. In addition to Instant3D (Li et al., 2023a) and MVDream (Shi et al., 2023b) as direct text-to-multi-view (T2MV) methods, we also adopt edge-cutting single image to multi-view (I2MV) methods CRM (Wang et al., 2024b), SV3D (Voleti et al., 2024) and Zero123++(Shi et al., 2023a). For these methods, we condition the diffusion model on the single view image generated by PixArt (prompted to generate CAD-style single object-centric image). The result of the CLIP score is 3 times averaged with different seeds. For FID, we use 3 different seeds for each of the 3K prompts to generate 9K images to test the distance with GT high-quality images.
@@ 4.2 Evaluation of Multi-view Images
Table 1: Benchmark of CLIP and FID score of text-to-multi-view (T2MV) models on generated 4 view images, CLIP score tests on 110 text prompts from GPTeval3D (Wu et al., 2024) while FID is measured with the distribution of 30K object-centric images generated by SOTA T2I models. For text-to-image-to-multi-view(T2I2MV), we input I2MV models with single view images generated by Pixart-$\alpha$, which superior single view image quality is marked in green.
Domain Method CLIP-R Score $\uparrow$ CLIP Score $\uparrow$ FID $\downarrow$
@@
T2I PixArt-$\alpha$ 96.1 94.7 25.9 41.5 20.7 5.4
T2I2MV SV3D 78.8 81.3 24.7 37.3 55.7 54.2
CRM 77.5 85.1 24.9 38.9 59.0 52.2
Zero123++ 78.0 84.5 24.2 36.9 53.2 49.3
T2MV Instant3D (unofficial) 83.6 91.1 25.6 39.2 83.2 77.9
MVDream 84.8 89.3 25.5 38.4 60.2 59.2
@@
Figure 6: Bootstrap3D can generate high quality multi-view images in out of domain cases compare to other edge-cutting multi-view diffusion models trained on Objaverse only.
As illustrated in Tab.1, compared to other methods, the T2MV diffusion model trained by our framework yields the best results both according to image-text alignment and image quality. For qualitative experiments, we visualize some of the comparisons with other edge-cutting multi-view diffusion model in Fig.6. For these image-to-multi-view models, we condition them on the top-left image generated by Bootstrap3D. Compared to these models trained solely on Objaverse Deitke et al. (2023), our model demonstrates superior generalizability when the image domain is beyond the domain of Objaverse. Since it is difficult to directly measure view consistency as there is no ground truth 3D object for text-to-3D generation. we evaluate the view consistency by synthesizing 3D objects through large reconstruction model in the following experiments. Qualitative results of real user cases are in Sup. A.1.
@@ 4.3 Evaluation of Generated 3D Objects
Table 2: Benchmark of CLIP and FID score of generated 3D objects based on rendered 9 view images. *MVDream is tested on 200 generated objects for FID test using SDS (Shi et al., 2023b), other methods are tested on 1000 objects using GRM (Xu et al., 2024b) and InstantMesh (Xu et al., 2024a) as sparse view reconstruction model.
Reconstruction Method CLIP-R Score $\uparrow$ CLIP Score $\uparrow$ FID $\downarrow$
CLIP-L/14 CLIP-bigG CLIP-L/14 CLIP-bigG PG2.5 PixArt
SDS MVDream* 85.2 90.8 26.1 39.4 57.8 56.7
GRM Instant3D (unofficial) 81.7 89.4 24.8 37.1 85.4 80.3
SV3D 74.1 82.8 23.4 34.1 68.4 69.1
Zero123++ 71.2 80.3 22.3 34.5 69.3 72.4
@@
InstantMesh Zero123++ 73.2 84.1 23.0 37.2 82.3 88.8
Bootstrap3D 87.1 92.0 26.0 39.2 61.2 55.3
View consistency is another crucial factor in reconstructing reasonable 3D objects. Miss alignment between different views can lead to blurred areas in reconstructed objects by large reconstruction model (Hong et al., 2023; Wei et al., 2024). This misalignment causes a significant deterioration in quality, resulting in a notable increase in metrics like FID. To assess the view consistency directly on 3D object, we employ GRM (Xu et al., 2024b) and InstantMesh (Xu et al., 2024a) to reconstruct the object given sparse view images generated in Sec. 4.2. We render 9 view images evenly in orbit for each object and evaluate the image-text alignment and image quality. As reported in Tab. 2. Bootstrap3D, after conditioning GRM or InstantMesh on 4 view images, can generate the best 3D objects both according to image-text alignment and image quality. GPT-4V based human-aligned evaluation based on GPTeval3D (Wu et al., 2024) is supplied in Sup. A.6.
We also present visualizations of some results in Fig.5. Bootstrap3D can generate objects with higher quality and prompt following ability. For other methods, as shown in the first column of Fig.5, although the first image may be well aligned with the given text prompt, the final 3D object may be compromised due to the limitations of its poor generalizability as they are also fine-tuned on Objaverse (Deitke et al., 2023) only. More visualizations and discussions of this are in Sup. A.2
@@ 4.4 Ablation Study
Table 3: Ablation study of proposed components and quantity of synthetic data. with CLIP-R Score represents image-text alignment and FID represents image quality.
@@
Synthetic data and dense captioning are proposed in our work to achieve high-quality images and better image-text alignment. We ablate their effects and the importance of data quantity in Tab. 3. Direct use of synthetic data without Training Timestep Reschedule (TTR) can cause severe blurs and deformation in final outcome. With the help of TTR, the mixture of data can not only improve image-text alignment but also maintain view consistency. Replacing Cap3D (Luo et al., 2024)’s caption with MV-LLaVA’s dense descriptive caption further improves the model’s capability of following prompts faithfully. Improvement through increasing volume of data also proves the scalability of our framework.
@@ 5 Conclusion and Discussion
Limitations and future work. Despite its promise, our work still faces several unresolved challenges. Firstly, the multi-view diffusion model is only the first step of the 3D content creation pipeline. Sparse view reconstruction models also need improvement as most edge-cutting sparse view reconstruction models are also trained on Objaverse Deitke et al. (2023) only. Secondly, Although MLLMs can estimate general quality and view consistency, subtle view inconsistency is hard to detect until ambiguity leads to blurred areas in reconstructed 3D object. While the proposed Training Timestep Reschedule can mitigate this problem, it cannot solve the problem fundamentally. Using synthetic data to train sparse view reconstruction models and quality estimation directly based on the reconstructed object are thus interesting future directions for improving 3D content creation.
Appendix A Appendix
A.1 Evaluation on wild prompts from real users
The results of the main part of the paper are only tested on GPT generated prompts. To test our work’s capability in wild cases, we also collect real user prompts and compare our method with Instant3D (Li et al., 2023a). specifically, we randomly collect 100 prompts from https://www.meshy.ai/ and test the CLIP-R precision as well as GPT based evaluation (detailed in Sup. A.6). Results and some qualitative cases are shown in Tab. 4 and Fig. 8. We highlight that our Bootstrap3D excels Instant3D (Li et al., 2023a) when tested on real user prompts through training on synthetic data.
Figure 8: Real user prompt cases visualization compared to Instant3D (Li et al., 2023a)
Table 4: Test results of in the wild cases. Bootstrap3D also excels Instant3D (Li et al., 2023a) in generating high quality images according to real user prompts.
@@ A.1 Evaluation on wild prompts from real users
Figure 9: Generated multivew images compare to other methods. Our method can generate multi-view images with long text control without encountering blurring effect from data generated by SV3D thanks to TTR and quality filtering.
A.2 More Visualization Compared to other Methods.
We show more visualization of the quantitative experiments shown in the main paper in Fig.9
For Image-to-3D methods, they can sometimes produces significant motion blurring and fails when the input image is out-of-distribution (like the 3rd cartoon style case). We resample the high-quality segment of the distribution of generated images using quality filtering based on MLLM methods. Furthermore, by employing TTR, we limit the impact of these data when training multi-view diffusion models, allowing our model to produce much clear results. In addition, we use a caption rewriting method, enabling finer prompt control for the generated multi-view images.
A.3 Data Statistics
A.3.1 Caption Analysis
Figure 10: Visualized analysis of dense reasoning descriptions generated by GPT4-Vision (OpenAI, 2023a) of the root noun-verb pairs (occurring over 1%) of the descriptions